# Covariance test and universal bootstrap by operator norm

### v1

@@ 5.1 Simulation
Table 1 reports the empirical sizes of the supremum, the Frobenius, and the operator norm tests at the significance level $0.05$ across various sample size $n$, dimension $p$, covariance $\bm{\Sigma}_{0}$ and data generation distributions of Gaussian and the uniform and t distributions, (distributions (1) and (2)). For results on Gaussian-uniform and Gaussian-$t$ distributions (distributions (3) and (4)), see Table 2 in the Supplement, which demonstrates similar patterns. The three operator norm tests are performed using the proposed universal bootstrap procedure, while the supremum and Frobenius norm tests rely on their respective asymptotic distribution formulas. Both Table 1 and Table 2 in the Supplement show that the universal bootstrap effectively controls the size of all operator norm-based statistics at the nominal significance level across all tested scenarios. This approach performs well under both Gaussian and non-Gaussian distributions, for both i.i.d. and non-i.i.d. data, in proportional and ultra-high dimensional settings, and across various covariance structures. These findings confirm the universality and consistency of the universal bootstrap procedure, providing empirical support for our theoretical results in Theorem 3.5 and Theorem 4.2. Additionally, the Frobenius norm-based test maintains appropriate size control, while the supremum norm-based test exhibits substantial size inflation at $n=100$ and moderate inflation even at a larger sample size of $n=300$.
To evaluate the empirical power of these statistics, we consider a setting where $\bm{\Sigma}=\bm{\Sigma}_{0}+\bm{\Delta}$, with $\bm{\Delta}$ having a low rank. This setup corresponds to the power analysis presented in Theorem 4.1. As shown in Theorem 4.1, the statistic $T$ outperforms $T^{\text{Roy}}$ when the eigenspaces of $\bm{\Delta}$ align with the eigenvectors of $\bm{\Sigma}_{0}$ corresponding to larger eigenvalues. In contrast, $T^{\text{Roy}}$ performs better when $\bm{\Delta}$ aligns with the smaller eigenvectors of $\bm{\Sigma}_{0}$. For a fair comparison, we define $\bm{\Delta}=\sigma\bm{u}\bm{u}^{T}/2+\sigma\bm{v}\bm{v}^{T}/4$, where $\bm{u}$ is the fifth-largest eigenvector of $\bm{\Sigma}_{0}$, $\bm{v}$ is uniformly sampled on the unit sphere in $\mathbb{R}^{p}$, and $\sigma$ is the signal strength. We term this configuration the spike setting. For completeness, we also conduct a power analysis where $\bm{\Delta}$ has full rank by setting $\bm{\Delta}=\sigma\bm{I}_{p}$, where $\sigma$ represents the signal level. This configuration, termed the white noise setting, represents the covariance structure after adding white noise to the null covariance. Together, the spike and white noise settings cover both low- and full-rank deviations from the null covariance. For illustration, we consider sample sizes $n=100$ and $300$, with dimension $p=1000$. Based on the universality results in Theorem 3.5 and empirical size performance, we conduct the power analysis using a Gaussian distribution for simplicity. Further empirical power comparisons under varying sample sizes and dimensions are provided in Table 1 of the Supplement table.
@@
Figure 1 illustrates the empirical power performance of six test statistics across different signal levels and null covariance structures under the spike setting, where the dashed line represents the significance level of $0.05$. Under this setting, the proposed operator norm-based statistics, $T$ and $T^{\text{Com}}$ exhibit similar power, while the two Frobenius norm-based statistics and the normalized operator norm statistic $T^{\text{Roy}}$ perform comparably. When $n=100$, the operator norm-based statistics $T$ and $T^{\text{Com}}$ maintain appropriate test size and achieve the highest power, approaching $1$ quickly across all three covariance configurations. The normalized operator norm statistic $T^{\text{Roy}}$ and the two Frobenius norm-based statistics $T^{\text{F},1}$ and $T^{\text{F},2}$ also control the test size but exhibit lower power, among which $T^{\text{Roy}}$ slightly outperforms. The supremum norm-based statistic, however, exhibits significant size inflation and fails to detect signals, as its power curve does not increase with signal level. The pattern is similar for $n=300$, except that the power curves for the supremum norm-based statistic remain near the dashed line, indicating limited power performance. This provides empirical evidence for the advantage of operator norm-based statistics over the other norms in the spike setting, particularly supporting the proposed statistic $T$ over $T^{\text{Roy}}$. We further make comparisons under the white noise setting across various signal levels in Figure 2. In this setting, the combined statistic $T^{\text{Com}}$ and the normalized statistic $T^{\text{Roy}}$ outperform other statistics. Because the white noise setting applies a uniform signal across eigenvector directions of $\bm{\Sigma}_{0}$ with both large and small eigenvalues, the normalized statistic $T^{\text{Roy}}$ achieves higher power than $T$, which is consistent with our power analysis Theorem 4.1. Notably, the combined statistic $T^{\text{Com}}$ performs robustly in both the spike and white noise settings, demonstrating enhanced power without size inflation, as supported by the generalized universal bootstrap consistency Theorem 4.2. As observed in the spike setting, Frobenius norm-based statistics display lower power, while the supremum norm-based statistic fails to control size. In summary, when the covariance structure is known, we recommend using the statistic $T$. When the covariance structure is unknown, we prefer the more robust combined statistic, $T^{\text{Com}}$.
@@ 5.2 Data application
We apply our test procedure to annual mean near-surface air temperature data from the global scale for the period $1960-2010$, using the HadCRUT4 dataset (Morice et al. (2012)) and the Coupled Model Intercomparison Project Phase 5 (CMIP5, Taylor, Stouffer and Meehl (2012)), as detailed in Li et al. (2021). The global dataset includes monthly anomalies of near-surface air temperature across $5^{\circ}\times 5^{\circ}$ grid boxes. To reduce dimensionality, these grid boxes are aggregated into larger $40^{\circ}\times 30^{\circ}$ boxes, resulting in $S=54$ spatial grid boxes. To mitigate distribution shifts due to long-term trends, we divide the $1960-2010$ period into five decades and analyze the data for each decade separately. The monthly data for each decade are averaged over five-year intervals, yielding a temporal dimension $T=2$ for each ten-year period. To illustrate the rationale of the covariance test in this context, we first outline the commonly used modeling procedure in climatology. Following Li et al. (2021), one considers a high-dimensional linear regression with errors-in-variables
@@
where $\bm{Y}_{k}\in\mathbb{R}^{L}$ is the vectorized observed mean temperature across spatial and temporal dimensions for the $k$-th decades after $1960$, with dimension $L=S\times T$. The vectors $\bm{X}_{k,1}$, $\bm{X}_{k,2},\bm{\epsilon}_{k}\in\mathbb{R}^{L}$ represent the expected unobserved climate response under the ANT and NAT forcings, and the unobserved noise, respectively. The coefficients $\beta_{k,1}$, $\beta_{k,2}$ are unknown scaling factors of interest for statistical inference. To estimate $\beta_{k,1}$, $\beta_{k,2}$, in addition to the observed data $\bm{Y}_{k}$, we have $n_{i}$ noisy observation fingerprints $\tilde{\bm{X}}_{k,i,j}$ of the climate system
@@
As noted by Olonscheck and Notz (2017), this equivalence (59) is crucial for optimal fingerprint estimation in climate models. However, few studies have validated (59) with statistical evidence, which calls for the need to test this hypothesis.
We construct the data matrix $\tilde{\bm{X}}_{k}\in\mathbb{R}^{n\times p}$ by combining quantities $\tilde{\bm{X}}_{k,1,j}-\bar{\tilde{\bm{X}}}_{k,1}$ for $j=1,\dots,n_{1}$ and $\tilde{\bm{X}}_{k,2,j}-\bar{\tilde{\bm{X}}}_{k,2}$ for $j=1,\dots,n_{2}$, where $n=n_{1}+n_{2}=81$ and $p=L=108$. Here $\bar{\tilde{\bm{X}}}_{k,i}=\sum_{j=1}^{n_{i}}\tilde{\bm{X}}_{k,i,j}/n_{i}$. We apply several statistics to the data matrix $\tilde{\bm{X}}_{k}$ with the hypothesized matrix $\bm{\Sigma}_{\bm{\epsilon},k}$: the proposed operator norm-based statistics $T$ (Opn) and $T^{\text{Com}}$ (Com), the operator norm-based statistics $T^{\text{Roy}}$ (Roy), the Frobenius norm-based statistics $T^{\text{F},1}$ (Lfn) and $T^{\text{F},2}$ (Ufn), and the supremum norm-based statistic $T^{\text{sup}}$ (Supn). For comparison, we generate i.i.d. Gaussian samples $\tilde{\bm{X}}_{k,j}^{\prime}\sim\mathcal{N}(\bm{0},\bm{\Sigma}_{\bm{\epsilon},k})$ for $j=1,\dots,n;\ k=1,\dots,5$ as a control group. The observed data $\tilde{\bm{X}}_{k}$ forms the observation group. The p-value results for these statistics are summarized in Table 2. The supremum norm-based statistic $T^{\text{sup}}$ fails to reject the observation group for all periods but rejects the control group in the periods $1960-1970$, $1970-1980$, and $1980-1990$. The Frobenius norm-based statistics $T^{\text{F},1}$ and $T^{\text{F},2}$ fail to reject the observation group in the periods $1970-1980$ and $1990-2000$, and $T^{\text{F},2}$ rejects the control group during $1990-2000$. Only the operator norm-based statistics $T$, $T^{\text{Com}}$, and $T^{\text{Roy}}$ reject the null hypothesis in the observation group while not rejecting it in the control group for all years. This suggests that the commonly assumed hypothesis (59) should be rejected, and a more suitable assumption should be used to estimate $\beta_{k,1}$ and $\beta_{k,2}$ for all $k$.
Table 2: Estimated p-values of tests for covariances for observation group and control group using the supremum norm-based statistic (Supn), the Frobenius norm-based statistic (Lfn, Ufn), and the operator norm-based statistics (Roy, Opn, Com), with sample size $n=81$, dimension $p=108$.
@@
In summary, the operator norm-based statistics demonstrate strong performance across various covariance structures and signal configurations, with the universal bootstrap ensuring the consistency of tests constructed from diverse combinations of operator norms. Numerical results align with the theoretical framework, indicating that the proposed tests exhibit robust power properties and that the universal bootstrap procedures maintain appropriate size control.
{funding}
Fang Yao is the corresponding author. This research is partially supported by the National Key Research and Development Program of China (No. 2022YFA1003801), the National Natural Science Foundation of China (No. 12292981, 11931001), the LMAM, the Fundamental Research Funds for the Central Universities, Peking University, and LMEQF.

### v2

@@ 5.1 Simulation
This approach performs well under both Gaussian and non-Gaussian distributions, for both i.i.d. and non-i.i.d. data, in proportional and ultra-high dimensional settings, and across various covariance structures. These findings confirm the consistency of the universal bootstrap procedure, providing empirical support for our theoretical results in Theorem 3.2 and Theorem 3.3. Additionally, the Frobenius norm-based test maintains appropriate size control, while the supremum norm-based test exhibits substantial size inflation at $n=100$ and moderate inflation even at a larger sample size of $n=300$.
To evaluate the empirical power of these statistics, we consider a setting where $\bm{\Sigma}=\bm{\Sigma}_{0}+\bm{\Delta}$, with $\bm{\Delta}$ having a low rank. This setup corresponds to the power analysis presented in Theorem 4.1. As shown in Theorem 4.1, the statistic $T$ outperforms $T^{\text{Roy}}$ when the eigenspaces of $\bm{\Delta}$ align with the eigenvectors of $\bm{\Sigma}_{0}$ corresponding to larger eigenvalues. In contrast, $T^{\text{Roy}}$ performs better when $\bm{\Delta}$ aligns with the smaller eigenvectors of $\bm{\Sigma}_{0}$. For a fair comparison, we define $\bm{\Delta}=\sigma\bm{u}\bm{u}^{T}/2+\sigma\bm{v}\bm{v}^{T}/4$, where $\bm{u}$ is the fifth-largest eigenvector of $\bm{\Sigma}_{0}$, $\bm{v}$ is uniformly sampled on the unit sphere in $\mathbb{R}^{p}$, and $\sigma$ is the signal strength. We term this configuration the spike setting. For completeness, we also conduct a power analysis where $\bm{\Delta}$ has full rank by setting $\bm{\Delta}=\sigma\bm{I}_{p}$, where $\sigma$ represents the signal level. This configuration, termed the white noise setting, represents the covariance structure after adding white noise to the null covariance. Together, the spike and white noise settings cover both low- and full-rank deviations from the null covariance. For illustration, we consider sample sizes $n=100$ and $300$, with dimension $p=1000$. Based on the universality results in Theorem 3.2 and empirical size performance, we conduct the power analysis using a Gaussian distribution for simplicity. Further empirical power comparisons under varying sample sizes and dimensions are provided in Table 1-4 of the Supplement table.
@@
Figure 1 illustrates the empirical power performance of six test statistics across different signal levels and null covariance structures under the spike setting, where the dashed line represents the significance level of $0.05$. Under this setting, the proposed operator norm-based statistics, $T$ and $T^{\text{Com}}$ exhibit similar power, while the two Frobenius norm-based statistics and the normalized operator norm statistic $T^{\text{Roy}}$ perform comparably. When $n=100$, the operator norm-based statistics $T$ and $T^{\text{Com}}$ maintain appropriate test size and achieve the highest power, approaching $1$ quickly across all three covariance configurations. The normalized operator norm statistic $T^{\text{Roy}}$ and the two Frobenius norm-based statistics $T^{\text{F},1}$ and $T^{\text{F},2}$ also control the test size but exhibit lower power, among which $T^{\text{Roy}}$ slightly outperforms. The supremum norm-based statistic, however, exhibits significant size inflation and fails to detect signals, as its power curve does not increase with signal level. The pattern is similar for $n=300$, except that the power curves for the supremum norm-based statistic remain near the dashed line, indicating limited power performance. This provides empirical evidence for the advantage of operator norm-based statistics over the other norms in the spike setting, supporting the proposed statistic $T$ over $T^{\text{Roy}}$.
We further make comparisons under the white noise setting across various signal levels in Figure 2. In this setting, the combined statistic $T^{\text{Com}}$ and the normalized statistic $T^{\text{Roy}}$ outperform other statistics. Because the white noise setting applies a uniform signal across eigenvector directions of $\bm{\Sigma}_{0}$ with both large and small eigenvalues, the normalized statistic $T^{\text{Roy}}$ achieves higher power than $T$, which is consistent with Theorem 4.1. Notably, the combined statistic $T^{\text{Com}}$ performs robustly in both the spike and white noise settings, demonstrating enhanced power without size inflation, as supported by the generalized universal bootstrap consistency Theorem 3.3. Similar to the spike setting, Frobenius norm-based statistics display lower power, while the supremum norm-based statistic fails to control size. In summary, when the covariance structure is known, we recommend using the statistic $T$. When the covariance structure is unknown, we prefer the more robust combined statistic, $T^{\text{Com}}$.
@@ 5.2 Data application
We apply our test procedure to annual mean near-surface air temperature data from the global scale for the period $1960-2010$, using the HadCRUT4 dataset (Morice et al., 2012) and the Coupled Model Intercomparison Project Phase 5 (CMIP5, Taylor et al., 2012), as detailed in Li et al. (2021) (https://figshare.com/articles/dataset/Uncertainty_in_Optimal_Fingerprinting_is_Underestimated/14981241/3?file=28924416). The global dataset includes monthly anomalies of near-surface air temperature across $5^{\circ}\times 5^{\circ}$ grid boxes. To reduce dimensionality, these grid boxes are aggregated into larger $40^{\circ}\times 30^{\circ}$ boxes, resulting in $S=54$ spatial grid boxes.
To mitigate distribution shifts due to long-term trends, we divide the $1960-2010$ period into five decades and analyze the data for each decade separately. The monthly data for each decade are averaged over five-year intervals, yielding a temporal dimension $T=2$ for each ten years. We first outline the commonly used modeling procedure in climatology. Following Li et al. (2021), one considers a high-dimensional linear regression
@@
where $\bm{Y}_{k}\in\mathbb{R}^{L}$ is the vectorized observed mean temperature across spatial and temporal dimensions for the $k$-th decades after $1960$, with dimension $L=S\times T$. The vectors $\bm{X}_{k,1}$, $\bm{X}_{k,2},\bm{\epsilon}_{k}\in\mathbb{R}^{L}$ represent the unobserved expected climate response under the ANT and NAT forcings, and the unobserved noise, respectively. The coefficients $\beta_{k,1}$, $\beta_{k,2}$ are unknown scaling factors of interest for statistical inference. To estimate $\beta_{k,1}$, $\beta_{k,2}$, in addition to the observed data $\bm{Y}_{k}$, we have $n_{i}$ noisy observation fingerprints $\tilde{\bm{X}}_{k,i,j}$ of the climate system
@@
As noted by Olonscheck & Notz (2017), this equivalence (28) is crucial for optimal fingerprint estimation in climate models. However, few studies have validated (28) with statistical evidence, which calls for the need to test this hypothesis.
Table 2: Estimated p-values for observation group and control group using the supremum norm-based statistic (Supn), the Frobenius norm-based statistic (Lfn, Ufn), and the operator norm-based statistics (Roy, Opn, Com), with sample size $n=81$, dimension $p=108$.
@@
We construct the data matrix $\tilde{\bm{X}}_{k}\in\mathbb{R}^{n\times p}$ by combining quantities $\tilde{\bm{X}}_{k,1,j}-\bar{\tilde{\bm{X}}}_{k,1}$ for $j=1,\dots,n_{1}$ and $\tilde{\bm{X}}_{k,2,j}-\bar{\tilde{\bm{X}}}_{k,2}$ for $j=1,\dots,n_{2}$, where $n=n_{1}+n_{2}=81$ and $p=L=108$. Here $\bar{\tilde{\bm{X}}}_{k,i}=\sum_{j=1}^{n_{i}}\tilde{\bm{X}}_{k,i,j}/n_{i}$. We apply several statistics to the data matrix $\tilde{\bm{X}}_{k}$ with the hypothesized matrix $\bm{\Sigma}_{\bm{\epsilon},k}$: the proposed operator norm-based statistics $T$ (Opn) and $T^{\text{Com}}$ (Com), the operator norm-based statistics $T^{\text{Roy}}$ (Roy), the Frobenius norm-based statistics $T^{\text{F},1}$ (Lfn) and $T^{\text{F},2}$ (Ufn), and the supremum norm-based statistic $T^{\text{sup}}$ (Supn). For comparison, we generate i.i.d. Gaussian samples $\tilde{\bm{X}}_{k,j}^{\prime}\sim\mathcal{N}(\bm{0},\bm{\Sigma}_{\bm{\epsilon},k})$ for $j=1,\dots,n;\ k=1,\dots,5$ as a control group. The observed data $\tilde{\bm{X}}_{k}$ forms the observation group. The p-value results for these statistics are summarized in Table 2. The supremum norm-based statistic $T^{\text{sup}}$ fails to reject the observation group for all periods but rejects the control group in the periods $1960-1970$, $1970-1980$, and $1980-1990$. The Frobenius norm-based statistics $T^{\text{F},1}$ and $T^{\text{F},2}$ fail to reject the observation group in the periods $1970-1980$ and $1990-2000$, and $T^{\text{F},2}$ rejects the control group during $1990-2000$. Only the operator norm-based statistics $T$, $T^{\text{Com}}$, and $T^{\text{Roy}}$ reject the null hypothesis in the observation group while not rejecting the control group for all years. This suggests that the commonly assumed hypothesis (28) should be rejected and replaced by a more suitable assumption to estimate $\beta_{k,1}$ and $\beta_{k,2}$.
SUPPLEMENTARY MATERIAL
Universal_Bootstrap_supp:
The additional theoretical results related to Section 3 and 4, more simulation examples and implementation details, all proofs and auxiliary lemmas are deferred to the supplementary material. (.pdf type)
Code and Data
Code to implement and reproduce the simulations and real data applications and corresponding output and raw datasets.